# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

### v3

@@ 5.3 The modified limit filter
Definition 5.48 (Modified limit filter).
Define the filter $\tilde{\mathscr{F}}^{*}_{\lambda}$ on $\mathscr{G}_{\lambda}$ inductively:
Stage 0: $\tilde{\mathscr{F}}^{*}_{0}$ is the filter on $\mathscr{G}_{0}=\operatorname{Sym}(\omega_{1})$ generated by $\{\operatorname{Fix}(E):E\subseteq\omega_{1}\text{ countable}\}$. Note that $\operatorname{Fix}(E)=\Delta_{0}^{\uparrow}(E,\varnothing)$, and $(E,\varnothing)$ is trivially globally admissible.
@@
    Pushforwards: $(\pi^{\alpha}_{\alpha+1})^{-1}[H]$ for $H\in\tilde{\mathscr{F}}^{*}_{\alpha}$;
@@
    $\Delta^{\uparrow}$-subgroups: $\Delta_{\alpha+1}^{\uparrow}(E,D)$ for countable $E\subseteq\omega_{1}$ and countable $D\subseteq\mathrm{Coords}_{<\alpha+1}$ such that $(E,D)$ is globally admissible at stage $\alpha+1$.
@@
    Pushforwards: $(\pi^{\beta}_{\lambda})^{-1}[H]$ for $\beta<\lambda$ and $H\in\tilde{\mathscr{F}}^{*}_{\beta}$;
@@
    $\Delta^{\uparrow}$-subgroups: $\Delta_{\lambda}^{\uparrow}(E,D)$ for countable $E\subseteq\omega_{1}$ and countable $D$ (package coordinates at stages $\lambda$) such that $(E,D)$ is globally admissible at stage $\lambda$.
@@
We require global admissibility in the $\Delta^{\uparrow}$-generator clause to ensure that the core lemma (Lemma 5.50) produces admissible pairs, which is needed for the properness proof (Lemma 5.52(iii)).
By Lemma 5.38, conjugation sends $\Delta_{\lambda}^{\uparrow}(E,D)$ to $\Delta_{\lambda}^{\uparrow}(g\cdot E,\,g\cdot D)$, and $g\cdot E$ and $g\cdot D$ remain countable (Lemma 5.30). The pair $(g\cdot E,\,g\cdot D)$ need not be globally admissible, but by Lemma 5.40 there is a countable $E^{\ast}\supseteq g\cdot E$ such that $(E^{\ast},\,g\cdot D)$ is globally admissible and $\Delta_{\lambda}^{\uparrow}(E^{\ast},\,g\cdot D)\leq\Delta_{\lambda}^{\uparrow}(g\cdot E,\,g\cdot D)$. Together with Corollary 5.39, this ensures that the normal filter generation process always contains $\Delta^{\uparrow}$-groups with globally admissible parameters.
@@
For every $H\in\tilde{\mathscr{F}}^{*}_{\lambda}$, there exist countable $E\subseteq\omega_{1}$ and countable $D$ (package coordinates at stages $\lambda$) with $(E,D)$ globally admissible such that:
@@
By definition of the normal filter generated by a family, every element $H\in\tilde{\mathscr{F}}^{*}_{\lambda}$ contains a finite intersection of conjugates of generators. We show that such intersections contain $\Delta^{\uparrow}$-groups with globally admissible parameters.
Conjugates of $\Delta^{\uparrow}$-generators: Let $\Delta_{\lambda}^{\uparrow}(E,D)$ be a globally admissible generator and let $g\in\mathscr{G}_{\lambda}$. By Lemma 5.38:
$g\cdot\Delta_{\lambda}^{\uparrow}(E,D)\cdot g^{-1}=\Delta_{\lambda}^{\uparrow}(g\cdot E,g\cdot D).$
Note that $g\cdot E$ and $g\cdot D$ are countable: $g\cdot E=\pi^{0}_{\lambda}(g)[E]$ is countable because bijections preserve cardinality, and similarly $g\cdot D$ is countable because the action on package coordinates is a bijection (Lemma 5.30(ii)). The parameters $(g\cdot E,g\cdot D)$ need not be globally admissible. However, by Lemma 5.40, there exists $E^{\ast}\supseteq g\cdot E$ such that $(E^{\ast},g\cdot D)$ is globally admissible and $\Delta_{\lambda}^{\uparrow}(E^{\ast},g\cdot D)\leq\Delta_{\lambda}^{\uparrow}(g\cdot E,g\cdot D)$.
Thus every conjugate of an admissible $\Delta^{\uparrow}$-generator contains a globally admissible $\Delta^{\uparrow}$-group.
Finite intersections: If $H_{1}\supseteq\Delta_{\lambda}^{\uparrow}(E_{1},D_{1})$ and $H_{2}\supseteq\Delta_{\lambda}^{\uparrow}(E_{2},D_{2})$ with $(E_{i},D_{i})$ globally admissible, then by Corollary 5.35:
$H_{1}\cap H_{2}\supseteq\Delta_{\lambda}^{\uparrow}(E_{1}\cup E_{2},D_{1}\cup D_{2}).$
By Corollary 5.39, $(E_{1}\cup E_{2},D_{1}\cup D_{2})$ is globally admissible.
Pullback generators: Let $H=(\pi^{\beta}_{\lambda})^{-1}[K]$ with $K\in\tilde{\mathscr{F}}^{*}_{\beta}$. By induction on $\beta$, there exist countable admissible $(E,D)$ with $D\subseteq\mathrm{Coords}_{\beta}$ and $\Delta_{\beta}^{\uparrow}(E,D)\leq K$. By Lemma 5.42, $\Delta_{\lambda}^{\uparrow}(E,D)\leq H$. The pair $(E,D)$ remains globally admissible at stage $\lambda$ (the admissibility condition $\operatorname{Fix}(E)\leq\operatorname{Stab}(d)$ depends only on how $d$ was introduced, not on which stage we evaluate it). ∎
@@
For $d\in\bigcup_{n}D_{n}$, choose $k$ with $d\in D_{k}$. Then $\operatorname{Fix}(\bigcup_{n}E_{n})\leq\operatorname{Fix}(E_{k})\leq\operatorname{Stab}(d)$, where the last inequality holds by global admissibility of $(E_{k},D_{k})$. ∎
@@
    $\tilde{\mathscr{F}}^{*}_{\lambda}$ is $\omega_{1}$-complete.
@@
Since $E$ is countable and $\omega_{1}$ is uncountable, there exist distinct $\beta,\gamma\in\omega_{1}\setminus E$. Let $\pi=(\beta\ \gamma)$ (transposition). Then $\pi\in\operatorname{Fix}(E)$.
By Lemma 5.43, the diagonal lift $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$ exists. By Lemma 5.44, $\widehat{\pi}^{\,D}\in\Delta_{\lambda}^{\uparrow}(E,D)$.
Since $\widehat{\pi}^{\,D}$ acts as $\pi$ on the Cohen base (Lemma 5.31(i)), it swaps coordinates $\beta$ and $\gamma$, so $\widehat{\pi}^{\,D}\neq\mathrm{id}$.
Hence $\Delta_{\lambda}^{\uparrow}(E,D)$ contains a nontrivial element, contradicting $\Delta_{\lambda}^{\uparrow}(E,D)\leq\{\mathrm{id}\}$.
(iv) $\omega_{1}$-completeness. Let $\{H_{n}:n<\omega\}\subseteq\tilde{\mathscr{F}}^{*}_{\lambda}$. By Lemma 5.50, for each $n$ there exists a globally admissible $(E_{n},D_{n})$ with $\Delta_{\lambda}^{\uparrow}(E_{n},D_{n})\leq H_{n}$.
By Corollary 5.35:
$\bigcap_{n<\omega}H_{n}\supseteq\Delta_{\lambda}^{\uparrow}\left(\bigcup_{n}E_{n},\bigcup_{n}D_{n}\right).$
By Corollary 5.51, $(\bigcup_{n}E_{n},\bigcup_{n}D_{n})$ is globally admissible. Since countable unions of countable sets are countable (in the metatheory, which has $\mathsf{AC}$), the right-hand side is a $\Delta^{\uparrow}$-generator of the filter.
Hence $\bigcap_{n}H_{n}\in\tilde{\mathscr{F}}^{*}_{\lambda}$. ∎
@@
    for every $(\sigma,p)\in\tau$, we have $\sigma\in\operatorname{HS}_{\lambda}$.
@@
For any $\mathbb{P}_{\lambda}$-name $\tau$:
$\operatorname{Stab}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}\iff\exists\text{ countable globally admissible }(E,D)\text{ such that }\Delta_{\lambda}^{\uparrow}(E,D)\subseteq\operatorname{Stab}(\tau).$
@@
$(\Leftarrow)$ By upward closure, since $\Delta_{\lambda}^{\uparrow}(E,D)\in\tilde{\mathscr{F}}^{*}_{\lambda}$ for globally admissible $(E,D)$. $(\Rightarrow)$ By Lemma 5.50. ∎
@@
By Lemma 5.52, the triple $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ is a symmetric system in the sense of [8, Ch. 15]. By the symmetric extension theorem [8, Lemma. 15.51], $M_{\lambda}\models\mathsf{ZF}$. ∎
@@
The final model $\mathcal{M}=\bigcup_{\alpha\in\mathrm{Ord}}M_{\alpha}$ satisfies $\mathsf{ZF}$.
@@
$\mathcal{M}$ is the directed union of the transitive models $\{M_{\alpha}:\alpha\in\mathrm{Ord}\}$, each satisfying $\mathsf{ZF}$ by Theorem 5.55.
For any $\mathsf{ZF}$ axiom instance with parameters $a_{1},\ldots,a_{n}\in\mathcal{M}$: choose $\lambda$ large enough that $a_{1},\ldots,a_{n}\in M_{\lambda}$. Apply the axiom in $M_{\lambda}$ to obtain the required witness in $M_{\lambda}\subseteq\mathcal{M}$.
The inclusion $M_{\beta}\subseteq M_{\lambda}$ for $\beta<\lambda$ is guaranteed by the pushforward generators: if $\tau\in\operatorname{HS}_{\beta}$ with $\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\beta}$, then $\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\supseteq(\pi^{\beta}_{\lambda})^{-1}[\operatorname{Stab}_{\mathscr{G}_{\beta}}(\tau)]\in\tilde{\mathscr{F}}^{*}_{\lambda}$, so $\tau\in\operatorname{HS}_{\lambda}$. ∎

### v4

@@ 5.3 The modified limit filter
Definition 5.50 (Modified stage filters).
Define by recursion on stages $\alpha$ a normal filter $\tilde{\mathscr{F}}^{*}_{\alpha}$ on $\mathscr{G}_{\alpha}$ as follows.
Stage 0: $\tilde{\mathscr{F}}^{*}_{0}$ is the normal filter on $\mathscr{G}_{0}=\mathscr{G}=\operatorname{Sym}(\omega_{1})$ generated by $\{\operatorname{Fix}(E):E\in[\omega_{1}]^{\leq\omega}\}$. (Note that $\operatorname{Fix}(E)=\Delta^{\uparrow}_{0}(E,\varnothing)$, and $(E,\varnothing)$ is trivially globally admissible.)
@@
    Pullbacks: $(\pi^{\alpha}_{\alpha+1})^{-1}[H]$ for $H\in\tilde{\mathscr{F}}^{*}_{\alpha}$;
@@
    $\Delta^{\uparrow}$-subgroups: $\Delta_{\alpha+1}^{\uparrow}(E,D)$ for countable $(E,D)$ globally admissible at stage $\alpha+1$.
@@
    Pullbacks: $(\pi^{\beta}_{\lambda})^{-1}[H]$ for $\beta<\lambda$ and $H\in\tilde{\mathscr{F}}^{*}_{\beta}$;
@@
    $\Delta^{\uparrow}$-subgroups: $\Delta_{\lambda}^{\uparrow}(E,D)$ for countable $(E,D)$ globally admissible at stage $\lambda$.
@@
We restrict the $\Delta^{\uparrow}$-generator clause in Definition 5.50 to globally admissible pairs so that later arguments can uniformly recover $\Delta^{\uparrow}$-subgroups with admissible parameters (notably Lemma 5.52, used in Lemma 5.54(iii),(iv)).
Normality interacts with admissibility as follows. If $(E,D)$ is globally admissible at stage $\lambda$ and $k\in\mathscr{G}_{\lambda}$, then Lemma 5.40 gives
$k\,\Delta_{\lambda}^{\uparrow}(E,D)\,k^{-1}=\Delta_{\lambda}^{\uparrow}(k\cdot E,\ k\cdot D),$
where $k\cdot E=\pi^{0}_{\lambda}(k)[E]$ and $k\cdot D$ is the induced action on coordinates (Lemma 5.32). The pair $(k\cdot E,\ k\cdot D)$ need not be globally admissible, but by Lemma 5.42 there is a countable $E^{\ast}\supseteq k\cdot E$ such that $(E^{\ast},\ k\cdot D)$ is globally admissible and
$\Delta_{\lambda}^{\uparrow}(E^{\ast},\ k\cdot D)\ \leq\ \Delta_{\lambda}^{\uparrow}(k\cdot E,\ k\cdot D).$
Together with Corollary 5.41 (and its countable variant Corollary 5.53), this ensures that the normal-filter generation process always contains $\Delta^{\uparrow}$-groups with globally admissible parameters.
@@
For each stage $\lambda$ and every $H\in\tilde{\mathscr{F}}^{*}_{\lambda}$, there exist countable $E\subseteq\omega_{1}$ and a countable $D\subseteq\mathrm{Coords}_{<\lambda}$ such that $(E,D)$ is globally admissible at stage $\lambda$ and
@@
We argue by induction on the stage $\lambda$.
Stage 0. By Definition 5.50, $\tilde{\mathscr{F}}^{*}_{0}$ is the normal filter on $\mathscr{G}_{0}=\operatorname{Sym}(\omega_{1})$ generated by $\{\operatorname{Fix}(E):E\in[\omega_{1}]^{\leq\omega}\}$. Every conjugate of $\operatorname{Fix}(E)$ is again of the form $\operatorname{Fix}(E^{\prime})$ for some countable $E^{\prime}$, and finite intersections satisfy $\operatorname{Fix}(E_{1})\cap\operatorname{Fix}(E_{2})\supseteq\operatorname{Fix}(E_{1}\cup E_{2})$. Hence every $H\in\tilde{\mathscr{F}}^{*}_{0}$ contains $\operatorname{Fix}(E)=\Delta^{\uparrow}_{0}(E,\varnothing)$ for some countable $E$, and $(E,\varnothing)$ is globally admissible. This proves the claim at stage 0.
Inductive step. Fix $\lambda>0$ and assume the statement holds for every $\beta<\lambda$. Let $H\in\tilde{\mathscr{F}}^{*}_{\lambda}$. By definition of “normal filter generated by a family”, there exist generators $G_{0},\dots,G_{n-1}$ from Definition 5.50 and elements $k_{0},\dots,k_{n-1}\in\mathscr{G}_{\lambda}$ such that
$\bigcap_{i<n}k_{i}G_{i}k_{i}^{-1}\ \subseteq\ H.$
We show that each conjugate $k_{i}G_{i}k_{i}^{-1}$ contains a globally admissible $\Delta^{\uparrow}_{\lambda}$-group.
Case 1: $G_{i}=\Delta_{\lambda}^{\uparrow}(E,D)$ with $(E,D)$ globally admissible at stage $\lambda$. By Lemma 5.40,
$k_{i}\,\Delta_{\lambda}^{\uparrow}(E,D)\,k_{i}^{-1}=\Delta_{\lambda}^{\uparrow}(k_{i}\cdot E,\ k_{i}\cdot D).$
Apply Lemma 5.42 to the countable set $k_{i}\cdot D$ and the countable set $k_{i}\cdot E$ to obtain a countable $E_{i}^{\ast}\supseteq k_{i}\cdot E$ such that $(E_{i}^{\ast},\ k_{i}\cdot D)$ is globally admissible. Then
$\Delta_{\lambda}^{\uparrow}(E_{i}^{\ast},\ k_{i}\cdot D)\ \leq\ \Delta_{\lambda}^{\uparrow}(k_{i}\cdot E,\ k_{i}\cdot D)=k_{i}G_{i}k_{i}^{-1}.$
Set $(E_{i},D_{i}):=(E_{i}^{\ast},\ k_{i}\cdot D)$.
Case 2: $G_{i}$ is a pullback generator. Then $G_{i}=(\pi^{\beta}_{\lambda})^{-1}[K]$ for some $\beta<\lambda$ and some $K\in\tilde{\mathscr{F}}^{*}_{\beta}$ (as in Definition 5.50). Put $\sigma:=\pi^{\beta}_{\lambda}(k_{i})\in\mathscr{G}_{\beta}$. Since $\pi^{\beta}_{\lambda}$ is a group homomorphism (Remark 5.22), one has
$k_{i}\,(\pi^{\beta}_{\lambda})^{-1}[K]\,k_{i}^{-1}=(\pi^{\beta}_{\lambda})^{-1}\!\big[\sigma K\sigma^{-1}\big].$
Because $\tilde{\mathscr{F}}^{*}_{\beta}$ is normal, $\sigma K\sigma^{-1}\in\tilde{\mathscr{F}}^{*}_{\beta}$. By the inductive hypothesis at stage $\beta$, choose a globally admissible pair $(E,D)$ at stage $\beta$ with $\Delta_{\beta}^{\uparrow}(E,D)\leq\sigma K\sigma^{-1}$. Then Lemma 5.44 yields
$\Delta_{\lambda}^{\uparrow}(E,D)\ \leq\ (\pi^{\beta}_{\lambda})^{-1}\!\big[\Delta_{\beta}^{\uparrow}(E,D)\big]\ \leq\ (\pi^{\beta}_{\lambda})^{-1}\!\big[\sigma K\sigma^{-1}\big]=k_{i}G_{i}k_{i}^{-1}.$
Since $\beta<\lambda$, the same pair $(E,D)$ is globally admissible at stage $\lambda$ as well. Set $(E_{i},D_{i}):=(E,D)$.
Assembling the finite intersection. Let $E^{\ast}:=\bigcup_{i<n}E_{i}$ and $D^{\ast}:=\bigcup_{i<n}D_{i}$. Since $n$ is finite, both unions are countable, and by iterating Corollary 5.41 the pair $(E^{\ast},D^{\ast})$ is globally admissible at stage $\lambda$. Moreover, by Corollary 5.37 (finite case),
$\Delta_{\lambda}^{\uparrow}(E^{\ast},D^{\ast})\ \leq\ \bigcap_{i<n}\Delta_{\lambda}^{\uparrow}(E_{i},D_{i})\ \leq\ \bigcap_{i<n}k_{i}G_{i}k_{i}^{-1}\ \subseteq\ H.$
This completes the induction and proves the lemma. ∎
@@
Put $E:=\bigcup_{n<\omega}E_{n}$ and $D:=\bigcup_{n<\omega}D_{n}$. Since we are working in the ambient universe in which the iteration is constructed, countable unions of countable sets are countable; hence $E\in[\omega_{1}]^{\leq\omega}$ and $D$ is countable. (No claim is made that such unions are countable inside the eventual symmetric model.)
To verify clause (iii) of Definition 5.28, fix $d\in D$ and choose $k$ with $d\in D_{k}$. Then $\operatorname{Fix}(E)\leq\operatorname{Fix}(E_{k})\leq\operatorname{Stab}(d)$, where the last inequality holds because $(E_{k},D_{k})$ is globally admissible. Therefore $(E,D)$ is globally admissible. ∎
@@
    $\tilde{\mathscr{F}}^{*}_{\lambda}$ is $\omega_{1}$-complete (i.e. closed under countable intersections).
@@
Since $E$ is countable and $\omega_{1}$ is uncountable, choose distinct $\beta,\gamma\in\omega_{1}\setminus E$ and let $\pi=(\beta\ \gamma)$. Then $\pi\in\operatorname{Fix}(E)$.
By Lemma 5.45, the diagonal lift $\widehat{\pi}^{\,D}\in\mathscr{G}_{\lambda}$ exists, and by Lemma 5.46 we have $\widehat{\pi}^{\,D}\in\Delta_{\lambda}^{\uparrow}(E,D)$. By Corollary 5.34, $\widehat{\pi}^{\,D}$ swaps $c_{\beta}\leftrightarrow c_{\gamma}$, so $\widehat{\pi}^{\,D}\neq\mathrm{id}$, contradicting $\Delta_{\lambda}^{\uparrow}(E,D)\leq\{\mathrm{id}\}$.
(iv) Let $\{H_{n}:n<\omega\}\subseteq\tilde{\mathscr{F}}^{*}_{\lambda}$. By Lemma 5.52, for each $n$ choose a globally admissible $(E_{n},D_{n})$ with $\Delta_{\lambda}^{\uparrow}(E_{n},D_{n})\leq H_{n}$. By Corollary 5.37,
$\bigcap_{n<\omega}H_{n}\ \supseteq\ \Delta_{\lambda}^{\uparrow}\Bigl(\bigcup_{n}E_{n},\ \bigcup_{n}D_{n}\Bigr).$
By Corollary 5.53, the pair $\bigl(\bigcup_{n}E_{n},\bigcup_{n}D_{n}\bigr)$ is globally admissible, and in the ambient universe where the iteration and filters are defined these unions are countable. Hence the right-hand side is one of the $\Delta^{\uparrow}$-generators in Definition 5.50, so by upward closure $\bigcap_{n<\omega}H_{n}\in\tilde{\mathscr{F}}^{*}_{\lambda}$. ∎
@@
    for every $\langle\sigma,p\rangle\in\tau$, we have $\sigma\in\operatorname{HS}_{\lambda}$.
@@
For any $\mathbb{P}_{\lambda}$-name $\tau$,
$\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}\iff\exists(E,D)\ \bigl[(E,D)\text{ is globally admissible at stage }\lambda\ \wedge\ \Delta_{\lambda}^{\uparrow}(E,D)\subseteq\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\bigr].$
@@
$(\Leftarrow)$ If $\Delta_{\lambda}^{\uparrow}(E,D)\subseteq\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)$ and $(E,D)$ is globally admissible, then $\Delta_{\lambda}^{\uparrow}(E,D)\in\tilde{\mathscr{F}}^{*}_{\lambda}$ (Definition 5.50), so by upward closure $\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}$.
$(\Rightarrow)$ Apply Lemma 5.52 to $H=\operatorname{Stab}_{\mathscr{G}_{\lambda}}(\tau)\in\tilde{\mathscr{F}}^{*}_{\lambda}$. ∎
@@
By Lemma 5.54(i),(ii), the triple $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ forms a symmetric system in the sense of [8, Ch. 15]. Therefore the associated symmetric extension $M_{\lambda}=\operatorname{HS}_{\lambda}^{G_{\lambda}}$ satisfies $\mathsf{ZF}$ by the symmetric extension theorem [8, Lemma 15.51]. ∎
@@
Let $G_{\infty}\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic and $G_{\lambda}:=G_{\infty}\cap\mathbb{P}_{\lambda}$. Let $\operatorname{HS}_{\mathrm{Ord}}$ be the class of hereditarily symmetric (cf. Definition 5.55) $\mathbb{P}_{\mathrm{Ord}}$-names (class-names) with respect to $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$, and define
$\mathcal{M}\ :=\ \{\dot{x}^{G_{\infty}}:\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}\}.$
Then $\mathcal{M}$ is a transitive class and $\mathcal{M}\models\mathsf{ZF}$.
@@
Work in the background metatheory used to formalize class-length iterations (cf. [5, §(Class-length iterations)] and Remark 5.59). In this setting, the forcing theorem holds for the definable class forcing $\mathbb{P}_{\mathrm{Ord}}$, and the class system $(\mathbb{P}_{\mathrm{Ord}},\mathscr{G}_{\mathrm{Ord}},\tilde{\mathscr{F}}^{*}_{\mathrm{Ord}})$ supports the usual recursion defining hereditarily symmetric class-names.
For each set ordinal $\lambda$, the restriction to stage $\lambda$ yields the set-length symmetric system $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ and hence $M_{\lambda}\models\mathsf{ZF}$ by Theorem 5.57. Moreover, every set-sized $\mathbb{P}_{\mathrm{Ord}}$-name $\dot{x}$ is supported in some stage $\lambda$, so $\dot{x}^{G_{\infty}}=\dot{x}^{G_{\lambda}}\in M_{\lambda}$. Consequently,
$\mathcal{M}=\{\dot{x}^{G_{\infty}}:\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}\}=\bigcup_{\lambda\in\mathrm{Ord}}M_{\lambda},$
so $\mathcal{M}$ is a transitive class.
The verification of the $\mathsf{ZF}$ axioms in $\mathcal{M}$ follows by the standard class-length symmetric iteration argument (ibid.): each axiom instance with set parameters from $\mathcal{M}$ is witnessed at some stage $\lambda$ and hence holds in $M_{\lambda}$, and therefore holds in $\mathcal{M}$. ∎